### v1

# Learning to Be A Doctor: Searching for Effective Medical Agent Architectures

Abstract
Large Language Model (LLM)-based agents have demonstrated strong capabilities across a wide range of tasks, and their application in the medical domain holds particular promise due to the demand for high generalizability and reliance on interdisciplinary knowledge. However, existing medical agent systems often rely on static, manually crafted workflows that lack the flexibility to accommodate diverse diagnostic requirements and adapt to emerging clinical scenarios. Motivated by the success of automated machine learning (AutoML), this paper introduces a novel framework for the automated design of medical agent architectures. Specifically, we define a hierarchical and expressive agent search space that enables dynamic workflow adaptation through structured modifications at the node, structural, and framework levels. Our framework conceptualizes medical agents as graph-based architectures composed of diverse, functional node types and supports iterative self-improvement guided by diagnostic feedback. Experimental results on skin disease diagnosis tasks demonstrate that the proposed method effectively evolves workflow structures and significantly enhances diagnostic accuracy over time. This work represents the first fully automated framework for medical agent architecture design and offers a scalable, adaptable foundation for deploying intelligent agents in real-world clinical environments.
Learning to Be A Doctor:
Searching for Effective Medical Agent Architectures
Yangyang Zhuang1, Wenjia Jiang1,2, Jiayu Zhang3,4, Ze Yang5 Joey Tianyi Zhou6,7 Chi Zhang1, 1AGI Lab, Westlake University, 2Henan University, 3Affiliated Hospital of Xuzhou Medical University 4Xuzhou Medical University 5Nanyang Technological University 6IHPC, Agency for Science, Technology and Research, Singapore 7CFAR, Agency for Science, Technology and Research, Singapore yyzhuang0211@gmail.com chizhang@westlake.edu.cn,
Figure 1: Workflow Evolution Diagram. This figure presents an iterative framework for workflow evolution in medical diagnosis tasks. Beginning with a structured input prompt that specifies diagnostic requirements, the system conducts process perception to analyze the current workflow’s structure and output behavior. Based on this analysis, both structural and prompt-level suggestions are generated to guide workflow refinement. These suggestions result in actionable modifications (e.g., adding, removing), which are subsequently integrated into the workflow through an iterative process. Through repeated iterations, the workflow is progressively optimized to improve diagnostic accuracy, logical coherence, and alignment with task objectives.
1 Introduction
Large Language Models (LLMs) have shown remarkable capabilities in solving a wide range of tasks. Leveraging these models, LLM-based agents are increasingly transforming how tasks are executed across various domains, such as automation, biology, and education. These agents are highly adaptive, capable of learning from prior experience, and continuously refining their behavior over time.
One particularly promising application of LLM-based agents is in the healthcare domain. The medical field demands high accuracy, efficiency, and scalability—traits that align well with the strengths of intelligent agents. In this setting, LLM-based agents are beginning to play crucial roles beyond serving as assistive tools. They are involved in essential processes such as diagnosis, medical data analysis, and treatment recommendation. By reducing the burden on healthcare professionals and improving clinical outcomes, these agents bring significant value to modern healthcare systems.
Among various agent system designs, multi-agent systems offer a natural fit for complex medical workflows. Medical imaging and diagnosis tasks generally involve multiple interdisciplinary departments, making coordination and collaboration essential. In this context, multi-agent systems can efficiently assign duties to corresponding specialized agents while maintaining coherence in overall decision-making. They also help bridge communication gaps between departments and enable standardized procedures across institutions.
However, current multi-agent systems face a key limitation in workflow design, relying on static, hand-crafted architectures tailored by domain experts. While effective in limited scenarios, these fixed workflows lack adaptability to evolving diagnostic techniques or new imaging modalities, requiring time-consuming redesigns. Moreover, for each new task or application scenario, developers are required to create customized workflows from scratch, limiting scalability and slowing down deployment.
To address the aforementioned challenges, we draw inspiration from the field of Automated Machine Learning (AutoML), particularly neural architecture search (NAS), to explore automated architecture design for medical agents. NAS has demonstrated its effectiveness in autonomously discovering high-performing neural architectures across a variety of tasks. Analogously, we argue that the automatic design of agent architectures can enhance the adaptability of multi-agent systems in dynamic clinical environments.
To this end, we conceptualize the agent system as a dynamic, graph-based workflow that evolves based on feedback from large language models (LLMs). To facilitate effective workflow evolution, we define a comprehensive search space comprising all valid evolution operations. These operations are categorized into three levels: node-level, structural-level, and framework-level. Node-level operations include adding or removing nodes and modifying attributes within individual nodes. To enable more sophisticated reasoning, structural-level operations introduce mechanisms such as conditional decision-making, iterative loops, and parallel execution, thereby enriching the logical workflow and enhancing diagnostic capabilities. At a higher abstraction level, framework-level operations incorporate advanced paradigms such as Chain of Thought Wei et al. (2022), Reflexion Shinn et al. (2023), and Roundtable Discussion Chen et al. (2024b), further improving the system’s reasoning and coordination abilities.
After defining evolution operations, we first established a naive feedforward workflow with only a single node to interact with the LLM and obtain the diagnostic results. As shown in Fig. 1, when diagnostic errors occur, the agent system can identify the root causes and provide both structural and prompt suggestions for workflow evolution. As more cases are processed, the workflow undergoes iterative refinement, progressively incorporating new nodes, structural logic, and framework-level reasoning strategies. Over time, it leads to the emergence of a more accurate, robust, and efficient diagnostic system. This continuous evolution strategy equips the agent system with the capacity for self-improvement to handle ever-changing real-world complexities.
To validate the effectiveness of our approach, we conduct comprehensive experiments on two medical diagnostic benchmarks. The results demonstrate that our method can automatically generate more efficient and accurate agent architectures, significantly outperforming strong baselines.
Our main contributions are as follows:
  • We propose the first framework for fully automated design of medical multi-agent systems using LLMs.
  • We introduce a novel hierarchical search space tailored for dynamic agent workflow evolution.
  • We develop a self-improving architecture search algorithm that enables agents to optimize their workflows through diagnostic feedback.
2 Related Work
2.1 LLM-based Agent in Medical Field
In recent years, the reasoning capabilities of large language models (LLMs) have improved significantly Chen et al. (2025), leading to their increasing integration into the medical domain to enhance both diagnostic accuracy and treatment effectiveness Hager et al. (2024); Schmidgall et al. (2024); Singhal et al. (2023). For instance, Agent HospitalLi et al. (2025) leverages scenario simulation to improve agents’ medical knowledge and diagnostic accuracy, offering a virtual environment for training and evaluation. Building on similar principles, SkinGPT-4Zhou et al. (2024) applies vision-language models to refine dermatological diagnosis, combining visual and textual data for precise outcomes. Meanwhile, MMedAgentLi et al. (2024) integrates toolchains to enable multi-modal task execution, supporting robust performance across varied medical applications. Additionally, some workKim et al. (2024); Li et al. (2023) focus on multi-agent collaboration improving efficiency in complex medical workflows. Nevertheless, despite their strengths in specific task performance, current systems pay limited attention to the dynamics workflows, which are essential for adapting to the complexity and variability of real-world medical environments.
2.2 Agent Collaboration and Self-Improving
Recent studies emphasize the growing importance of agent collaboration and self-improvement in multi-agent systems (MAS) based on large language models (LLMs). Common architectures include peer-to-peer setups, centralized controllers, and hierarchical frameworks Chen et al. (2023); Tran et al. (2025). Collaborative and self-improving MAS have shown promise in diverse application domains, like MetaGPTHong et al. (2024), LLM-BlenderJiang et al. (2023) and Multi-Agent DebateLiang et al. (2024). Beyond these domains, recent work focuses on enhancing adaptability in open-ended settings by exploring self-evolving mechanisms: EvoMAC Hu et al. (2024) introduces a self-evolving collaboration network trained via textual backpropagation for software-level development tasks. AutoAgentsChen et al. (2024a) supports automatic agent generation and flexible collaboration. AFLOWZhang et al. (2025) uses graph search to optimize task execution paths.
3 Methodology
Figure 2: Illustration of node attributes and hierarchical search space. (a) Demonstrates the attributes of a basic node and how it can evolve into complex workflow structures such as conditional branches, loops, and parallel executions. Each basic node contains configurable parameters including system prompts, human prompts, descriptions, and optional image paths. These attributes allow the agent to execute specific medical diagnostic tasks within the workflow. (b) illustrates the proposed hierarchical search space for dynamic workflow design, consisting of node-level, structural-level, and framework-level operations. These enable fine-grained node manipulation, control flow structures (e.g., conditionals, loops, parallelism), and emergent high-level reasoning patterns such as Chain-of-Thought, Reflexion, and Round Table Discussion.
We propose a dynamic framework that leverages large language models to optimize medical workflows. In our approach, workflows are represented as graph-based structures comprising heterogeneous node types, each characterized by specific attributes and functional roles. We first introduce the graph-based representation of medical workflows, where each node encapsulates a distinct operation with semantic and procedural properties. We then detail how combinations of these nodes define a structured and constrained search space, effectively capturing the spectrum of valid workflow configurations. Finally, we present the Workflow Evolution process to enabling continuous improvement in workflow design and performance.
3.1 Workflow Node Definition
In this section, we define the concept of a medical workflow and its foundational components, which collectively form the basis for dynamic task execution within the proposed framework. A workflow is modeled as a sequence of interconnected nodes, categorized into basic and tool types. A basic node directly interacts with the LLM, generating a single result based on the provided prompts. In contrast, a tool node utilizes predefined tools accessible to the LLM, with an emphasis on supporting workflow evolution. For simplicity, only one tool, the image search tool, is introduced. This tool accepts an image as input and returns the top k most similar disease labels along with their similarity scores, which serve as diagnostic references for the LLM.
Basic nodes serve as the foundation from which other structures evolve; for example, Figure  2 (a) show that a basic node combined with a looping condition transforms into a loop node, while one with a conditional branch becomes a conditional node. Multiple basic nodes running concurrently create a parallel structure. Each structural type plays a distinct role in facilitating the flexible and efficient execution of complex medical workflows. For the basic node type, the following key attributes are included:
  • System Prompt: Defines the role or identity of the agent within the node.
  • Human Prompt: Specifies the instructions for the agent to carry out the designated task.
  • Description: Provides a concise explanation of the node’s function or purpose.
For our tool node type, the following key attributes are included:
  • Tool Name: The name of the tool the node will use.
  • Description: Provides a concise explanation of the node’s function or purpose.
3.2 Search Space
To facilitate systematic workflow evolution, we propose a hierarchical and expressive search space that defines actionable strategies for agents across node-level, structural-level, and framework-level operations, enabling precise adjustments to workflow architecture and task-specific parameters.
Node-Level Operations. Node-level operations focus on modifying individual workflow components to optimize functionality and streamline execution.
  • Add: Introduces a new node to the workflow, which can be a basic functional node or a tool node integrated with specialized capabilities. By adding nodes, workflows can accommodate additional capabilities, enhancing problem-solving capabilities.
  • Remove: Eliminates unnecessary or redundant nodes to improve workflow efficiency. Removing nodes helps streamline execution paths, resulting in a more concise and targeted process.
  • Modify: Adjusts prompt configurations for existing nodes. This enables fine-grained control over node behavior to better fulfill task requirements.
Structural-level Operations. Structural operations enable the modification and expansion of overarching workflow logic, addressing branching, feedback loops, and parallel execution paths.
  • Conditional Structures: Conditional branches redirect workflow execution based on specified criteria, creating dynamic paths for tailored task handling. This adds flexibility and adaptability to workflows by allowing different paths based on decision logic and runtime conditions.
  • Loop Structures: Feedback loops allow iterative node execution under specific conditions, supporting tasks that require repeated evaluation or refinement. Proper loop designs optimize iterative processes, including condition thresholds, iteration limits, and feedback paths.
  • Parallel Structures: Parallel execution paths enable concurrent multi-agent analysis, enhancing computational efficiency and fostering collaborative workflows. These structures are particularly beneficial when tasks can be divided into independent sub-tasks that can be processed simultaneously, thereby reducing the overall execution time. Situations that demand high throughput, involve large data sets, or require simultaneous execution of independent processes can greatly benefit from parallel structures. Additionally, parallel execution allows different agents to express their viewpoints independently without being influenced by others, promoting unbiased contributions. Once all agents have shared their perspectives, they can proceed to integrate and consider each other’s insights, leading to a more comprehensive and collaborative outcome.
Framework-Level Design. The search space implicitly includes high-level conceptual workflow patterns—such as roundtable discussions—as abstract templates. These patterns are not pre-defined as fixed modules, but rather emerge dynamically through the composition of lower-level nodes and structural operations during the search process, as shown in figure2. For instance, the behavior of a roundtable discussion can be simulated by orchestrating a set of interconnected nodes in a looped structure. This design allows the framework to flexibly construct complex task-specific workflows by synthesizing reusable primitives in a goal-driven manner.
  • Reflection Framework: This framework focuses on iterative self-assessment and knowledge refinement to boost problem-solving abilities through reflective thinking. Promoting continuous evaluation and learning allows agents to adapt and improve over time, leading to more informed decisions.
  • Collaboration Framework: By allowing experts from various fields to work concurrently, this framework enhances the ability to address complex problems with better solutions.
  • Chain-of-Thought Reasoning: It decomposes complex tasks into sequential, modular steps, allowing for detailed analysis at each stage. By systematically addressing each component, it enhances the accuracy and thoroughness of the overall execution process.
  • Round Table Discussion Framework: Multiple agents collaborate and deliberate around a common task or theme to achieve a more efficient or accurate solution. Each agent may employ distinct models or algorithms and share information with one another, akin to a round table meeting, to integrate diverse insights and optimize the final outcome.
  • Conquer-and-Merge Discussion (CMD) Framework: Parallel analytical groups with voting mechanisms for consensus driven decision making.
3.3 Workflow Evolution
Our proposed workflow evolution process is a continuous iterative cycle. Initially, large language models perform a diagnostic analysis of the workflow execution to identify potential issues, categorizing them into various error cases. The causes of these errors are collected and analyzed to generate improvement suggestions. These suggestions are then transformed into actionable workflow modifications, aligning with the requirements to ensure targeted enhancements. The updated workflow is validated and subsequently updated, after which the iteration continues with the next diagnostic phase. This process of diagnosing, collecting causes, analyzing, receiving suggestions, verifying, and updating ensures that the workflow is continually refined and optimized for both structure and execution performance.
First, LLM focuses on diagnosing workflow errors, and it aggregates outputs from all workflow nodes and classifies errors into two primary categories. Image Understanding Errors: Issues arising from inaccuracies in feature recognition or extraction, which compromise the accuracy of diagnostic decisions. Diagnostic Decision Errors: Mistakes emerging from flawed reasoning or incorrect decision-making, even when image features are correctly identified. Then, LLM traces the root causes of these errors, determining whether they originate from inaccurate image feature recognition or deficiencies in diagnostic reasoning. Critical data related to the problematic nodes is preserved during this process, enabling a more in-depth analysis to uncover the origins of diagnostic inaccuracies.
Based on this analysis, the system proceeds to analyze the workflow’s logical structure and extract detailed descriptive metadata from its components. This process is facilitated by the visual perception capabilities of the large language model, which enable the generation of a structured, graph-based representation of the workflow. The resulting representation provides both a visual and logical overview of the workflow architecture, clearly illustrating the interconnections and dependencies among components. The workflow is visualized using Mermaid, which allows for direct viewing of loop conditions and conditional branches within the generated flowchart. This structural representation allows the system to accurately identify the relationships between nodes and understand the specific functional roles they fulfill within the overall process. By integrating this structural understanding with an analysis of key issues, the LLM can offer more targeted and effective suggestions for improving the workflow.
Next these suggestions undergo a filtering process to eliminate impractical proposals, such as those requiring external inputs beyond the scope of the intelligent agent’s capabilities. The remaining suggestions are further categorized into two groups: Structural suggestions: Suggestions aimed at optimizing the overall workflow architecture and structural dependencies. Prompt suggestions: Suggestions focused on refining task-specific configurations, such as prompt templates, to enhance node performance.
Finally, these filtered suggestions are refined and integrated into the workflow to support effective evolution and optimization. Initially, the suggestions are reformulated into actionable updates and returned in standardized formats as JSON. The structural suggestions are primarily aimed at improving the overall architecture of the workflow. Due to their potential impact on the underlying structure, these suggestions must undergo a rigorous validation process. This step is essential to ensure that no invalid or inappropriate actions from the search space are introduced, as such actions could disrupt or compromise the integrity of the existing workflow design. Once validated, they are applied to optimize the workflow’s architecture and execution flow. Prompt suggestions, on the other hand, directly modify task-specific configurations, such as prompt templates, and are applied seamlessly without requiring additional validation steps. By systematically refining and integrating these suggestions, the workflow evolves iteratively, achieving continuous improvement in both its structure and diagnostic performance.
Through validation and integration, the framework systematically incorporates improvements while preventing the introduction of inefficient or erroneous actions that could disrupt the workflow structure. This refinement mechanism ensures that every validated structural recommendation and prompt modification contributes to meaningful enhancements in workflow functionality, enabling continuous and targeted optimization.
The entire process—from diagnosis and error analysis to workflow updates—can be viewed as an evolutionary cycle. During each iteration, we run multiple examples and then refine the workflow to produce an updated version. This process is repeated continuously until the final workflow’s accuracy on the validation set converges, signifying the completion of the evolution.
4 Experiments
4.1 Experimental Setup
Figure 3: Workflow evolution over iterations. We visualize the diagnostic workflow at three key stages: the initial state (Iteration 0), an intermediate state after partial refinement (Iteration 3), and the final evolved workflow (Iteration 6). The evolution process demonstrates how the system incrementally improves its structure by adding conditional branches, parallel reasoning paths, and error-specific loops. These structural refinements enable more accurate and robust diagnostic reasoning through iterative self-adaptation.
Datasets. To evaluate the effectiveness of our proposed framework, we select the task of skin disease diagnostics as the primary testbed. This domain is particularly suitable for evaluating multi-agent architectures due to its inherent complexity: accurate diagnosis often requires interdisciplinary knowledge and a multi-step reasoning process involving visual inspection of skin lesions, patient history, and symptom-diagnosis correlations. We utilized two dermatological datasets to evaluate the workflow evolution:
SKIN Concepts Dataset. SKINCONDaneshjou et al. (2023) builds upon the Fitzpatrick 17k skin disease datasetGroh et al. (2022, 2021), including 3230 images annotated with 48 clinically relevant concepts, such as "plaque," "scale," and "erosion." These annotations span 113 types of skin diseases and are based on established clinical vocabulary used by dermatologists, ensuring their applicability across diverse disease processes.
Augmented Skin Conditions Image Dataset. Skin Conditions Naqvi (2023)contains a total of 2,394 augmented images across six different dermatological conditions, with each category comprising 399 images. This balanced structure makes it ideal for training machine learning models, particularly in the field of medical image analysis.
Data Processing. During preliminary data preparation, we identified a subset of images within the datasets that depicted skin conditions located in private or sensitive body areas. Due to content moderation constraints, these images were rejected by the LLM and could not be processed during workflow execution. Additionally, we encountered another issue where approximately 20 images had broken URL links, preventing them from being downloaded. To maintain the integrity of our experiments and ensure consistent, reliable evaluation across all agent architectures, we excluded these samples from the dataset.
Datasets Split: SKIN Concepts Dataset: The training and validation sets contain 50 randomly sampled images each, and the test set includes 224 images, ensuring that all disease types are represented with two images per class. SKIN Conditions Dataset: Training and validation sets each have 50 images; the test set contains 120 images, with 20 per condition.
Implementation Details. We implemented an experimental process that incorporates training, validation, and testing phases.
After each training batch, the evolutionary mechanism adjusts the workflow structure. The model’s performance is then validated on a separate set. This cycle of training and evolution repeats until training completion, followed by a final evaluation of the test set.
Our framework is built on LangGraphInc. (2023). LangGraph provides us with some pre-packaged content, and we only need to develop our node designs on top of LangGraph. Additionally, we provide a tool list where you can add your custom tools. In the tool list, we have defined an image search tool for finding similar images. Image search tool utilizes the pre-trained CLIP(ViT-L/14)Radford et al. (2021) model to extract feature vectors without any further fine-tuning from images for data that is not part of the training, validation, and testing sets, storing these vectors in the PineconePinecone Systems (2023) database. When searching for similar images using the image tool search, it can directly match the feature vectors stored in this database.
Figure 4: Top-k diagnostic accuracy over workflow evolution iterations. This figure shows how the Top-k accuracy improves as the workflow undergoes iterative optimization. Each point represents the diagnostic performance after one round of workflow evolution. The increasing trend demonstrates that structural refinements and prompt updates effectively enhance diagnostic capability. The curve eventually converges, indicating that the system stabilizes once a sufficiently optimized workflow is discovered. The results in the figure were obtained using the GPT-4o.
Baselines. To establish a comparative framework for evaluating the effectiveness of our workflow evolution methodology, we implemented two baseline approaches for image-based dermatological diagnosis. The first baseline utilizes GPT-4o as a single-agent model to perform direct diagnosis of input images, without any intermediate reasoning or workflow adjustments. This provides a straightforward benchmark for assessing diagnostic accuracy.
The second baseline augments GPT-4o with a Chain-of-Thought (CoT) prompting mechanism. This approach guides GPT-4o to reason through intermediate steps before producing the final diagnosis. Although CoT enhances interpretability compared to direct diagnosis, it does not leverage adaptive or evolving workflows, which are central to our proposed methodology.
Figure 5: Per-disease consensus accuracy (cons@64) across different methods.This figure illustrates the prediction stability of each method under six different disease categories, evaluated via majority-vote consensus over 64 samples per class. Our method consistently achieves higher cons@64 scores, indicating improved diagnostic robustness and intra-class agreement across cases.
The third baseline involves a "round table" discussion setup, which is a type of multi-agent system configuration. In this approach, multiple agents collaborate and deliberate around a common task or theme to achieve a more efficient or accurate solution. Each agent may employ distinct models or algorithms and share information with one another, akin to a round table meeting, to integrate diverse insights and optimize the final outcome.
These baselines allow for quantifying the benefits of workflow evolution in terms of generalization ability and diagnostic performance.
LLM Method Skin Concepts Accuracy (%) Skin Conditions Accuracy (%) Top-1 Top-3 Top-5 Top-1 Top-3 Top-5 GPT-4o IO 20.27 30.63 36.04 50.83 78.33 86.67 CoT Wei et al. (2022) 18.47 28.83 33.78 55.83 76.67 82.50 Round Table Chen et al. (2024b) 21.17 27.93 32.43 45.83 75.83 80.83 Ours 29.28 40.09 50.45 90.83 95.00 100.00 GPT-4o-mini IO 11.71 20.72 23.87 27.50 69.17 80.83 CoT Wei et al. (2022) 6.31 15.32 24.32 22.50 65.00 84.17 Round Table Chen et al. (2024b) 10.81 19.82 23.42 25.83 70.00 78.33 Ours 13.51 21.62 24.77 45.83 74.17 85.83
Table 1: Top-k diagnostic accuracy (%) of different methods using GPT-4o and GPT-4o-mini on skin concepts and skin conditions. Higher values indicate better performance.
Operation Top-1 Acc. (%) Top-3 Acc. (%) Top-5 Acc. (%) Add Tool Node 21.62 (-7.66) 30.18 (-9.91) 36.94 (-13.51) Modify Node Prompt 19.37 (-9.91) 27.93 (-12.16) 33.78 (-16.67) Remove Node 28.83 (-0.45) 41.44 (+1.35) 50.90 (-0.45)
Table 2: Ablation results of different operations. Each row shows the Top-1, Top-3, and Top-5 accuracies when a specific operation is disabled. Values in parentheses indicate accuracy drop compared to the full model Top-1(29.28.%), Top-3(40.09%), and Top-5(50.45%).
Evaluation Metrics. To evaluate the diagnostic models, we adopted accuracy-based metrics across Top-1, Top-3, and Top-5 predictions. Each diagnostic attempt produces a ranked set of predictions based on the model’s confidence scores, and the correctness of these predictions is compared against ground truth labels.
During the validation phase, the Top-1, Top-3, and Top-5 accuracies are measured after each evolution iteration, enabling us to observe how these metrics improve as the workflow evolves. In the test phase, the final accuracies obtained with our workflow evolution method are compared against those of the baselines, demonstrating the advantages of dynamic and adaptive workflows.
Following Wang et al. (2023), we adopt the majority voting consensus metric, denoted as cons@64, which reflects how often a method arrives at a consistent diagnosis across multiple inputs of the same class (n=64). This metric offers a proxy for model robustness and reliability under input variance and sampling noise.
These metrics allow us to quantify both the progression of accuracy during workflow evolution and the overall generalization ability of the proposed method when tested on unseen data, particularly in diverse disease categories.
4.2 Analysis of Evolutionary Processes
Evolution Visualization. If the workflow nodes are directly input to the LLM in text form, the LLM may easily overlook structural relationships such as loops and conditional branches. To better understand how our workflow evolves over time, we visualize the structure at three representative stages: the initial workflow (Iteration 0), an intermediate state (Iteration 3), and the final evolved workflow (Iteration 6). These visualizations reveal how the system progressively incorporates structural refinements, such as conditional branches, parallel reasoning paths, and error-specific loops. Figure 3 presents a comparative view of these stages.
Quantitative Analysis To quantitatively evaluate the effectiveness of our workflow evolution framework, we track the diagnostic performance—measured by Top-1 accuracy, Top-3 accuracy and Top-5 accuracy—across multiple evolution iterations. At each iteration, the system refines its workflow based on prior diagnostic errors and structural improvements, allowing it to incrementally enhance its reasoning capabilities.
Figure 4 illustrates how the Top-1 accuracy, Top-3 accuracy, and Top-5 accuracy change as the workflow evolves. The straight line represents the accuracy of directly inputting images into the LLM for diagnosis. It can be observed that the accuracy of the evolved workflow is always higher than the IO line. We observe a consistent upward trend during the early iterations, indicating that the refinement of workflow structure and prompt configurations leads to improved diagnostic performance. The accuracy eventually converges, suggesting that the system has reached a stable and effective workflow configuration. This pattern demonstrates the framework’s ability to learn from experience and self-optimize over time.
Additionally, in the Skin Concepts dataset, both the training and validation sets consist of 50 randomly selected images, while the testing set includes images from all categories. Our workflow iteratively evolves on the training set and is validated on the validation set. Ultimately, it achieves good results on the testing set, which includes all categories. This demonstrates the generalization capability of our evolved workflow.
4.3 Performance Comparison Results
Diagnostic Accuracy Comparison. To comprehensively evaluate the effectiveness of our method in dermatological diagnostic tasks, we compare various prompting strategies across multiple state-of-the-art large language models (LLMs). Specifically, we report the Top-k diagnostic accuracy on both skin concepts and skin conditions using GPT-4o and GPT-4o-mini. As shown in Table 1, our approach consistently outperforms existing baselines, including IO, CoT Wei et al. (2022), and Round TableChen et al. (2024b), achieving superior results across all LLMs and evaluation metrics. As shown in Figure 1, our method achieves substantial gains in Top-k accuracy, highlighting its superior diagnostic capability and robustness across different models.
Workflow Stability. In addition to diagnostic accuracy and structural adaptability, we assess the stability of our evolved workflows by evaluating their consensus behavior across multiple independent samples. We report per-class consensus accuracy (cons@64) for each method using the GPT-4o model on the Skin Concepts dataset, as shown in Figure 5. Our method consistently achieves higher consensus scores across a wide range of conditions, including both common (e.g., eczema, psoriasis) and rare or visually complex diseases (e.g., impetigo, rosacea). This demonstrates not only the general robustness but also the fine-grained reliability of the evolved workflows.
4.4 Component Ablation Analysis
We conduct a comprehensive ablation study to investigate the role of different workflow modification actions in our evolution framework. Specifically, we categorize all actions into three main types: Add, Modify, and Remove. Each of these categories encompasses a set of fine-grained operations applied to either nodes or structural components (e.g., conditional, loop, parallel).
In this study, we conducted ablation experiments to evaluate the impact of different operations on the effectiveness of the workflow. Specifically, we selected three key operations from the set of possible actions: adding a tool node, modifying a node prompt, and removing a node. Firstly, the operation of adding a tool node is intended to acquire information about similar images. To assess the necessity of this operation, we removed it from the search space and observed the performance changes. Similarly, we applied the same treatment to the operations of modifying a node prompt and removing a node, excluding them from the search space to analyze each operation’s contribution to the overall system performance and effectiveness. These experiments will aid in understanding the role and significance of each operation in workflow optimization.
To assess their individual contribution, we disable one specific action at a time and examine the resulting Top-1 diagnostic accuracy on the Skin Concepts dataset. The results are summarized in Table 2, which reveals the relative importance of each operation in enabling effective performance improvements. On the one hand, notably, actions related to adding tool node and prompt modification exhibit a particularly strong impact on performance, underscoring their significance in workflow adaptability and optimization. On the other hand, we can see that removing nodes does not significantly impact accuracy. However, in practice, having too many nodes with similar functions can lead to decreased execution efficiency of the workflow.
5 Conclusion
Current dermatology-focused multi-agent systems are hindered by rigid, manually designed workflow frameworks that lack the adaptability to meet evolving diagnostic demands and diverse requirements. To overcome this limitation, we propose a node-level evolution framework that introduces flexibility through autonomous exploration and development of feasible workflows. By enabling dynamic evolution at the node level, our framework allows workflows to adapt organically, facilitating efficient responses to changing conditions and complex diagnostic scenarios. This approach significantly enhances flexibility by supporting modular updates, structural adjustments, and parallel processing while autonomously optimizing workflow paths. As a result, our framework offers a scalable and adaptive solution for improving diagnostic accuracy, operational efficiency, and responsiveness in dermatological multi-agent systems.

### v2

@@
 This figure presents an iterative framework for workflow evolution in medical diagnosis tasks.
-Beginning with a structured input prompt that specifies diagnostic requirements, the system conducts process perception to analyze the current workflow’s structure and output behavior.
+Starting with an initial workflow, the system performs process perception to analyze its structure and outputs.
 Based on this analysis, both structural and prompt-level suggestions are generated to guide workflow refinement.
-These suggestions result in actionable modifications (e.g., adding, removing), which are subsequently integrated into the workflow through an iterative process.
-Through repeated iterations, the workflow is progressively optimized to improve diagnostic accuracy, logical coherence, and alignment with task objectives.
+Through repeated iterations, the workflow is progressively optimized to improve diagnostic accuracy and alignment with task objectives.
 1 Introduction
-Large Language Models (LLMs) have shown remarkable capabilities in solving a wide range of tasks.
-Leveraging these models, LLM-based agents are increasingly transforming how tasks are executed across various domains, such as automation, biology, and education.
+Large Language Models (LLMs) Mann et al.
+(2020); Workshop et al.
+(2022); Achiam et al.
+(2023) have shown remarkable capabilities in solving a wide range of tasks Park et al.
+(2023); Shinn et al.
+(2023b).
+Leveraging these models, LLM-based agents are increasingly transforming how tasks are executed across various domains, such as automation Li et al.
+(2024b); Zhang et al.
+(2025a); Jiang et al.
+(2025) and biology Li et al.
+(2024a, 2025).
 These agents are highly adaptive, capable of learning from prior experience, and continuously refining their behavior over time.
-One particularly promising application of LLM-based agents is in the healthcare domain.
-The medical field demands high accuracy, efficiency, and scalability—traits that align well with the strengths of intelligent agents.
-In this setting, LLM-based agents are beginning to play crucial roles beyond serving as assistive tools.
+One particularly promising application of LLM-based agents Kim et al.
+(2024); Jiang et al.
+(2025) is in the healthcare domain Wang et al.
+(2024b); Li et al.
+(2023).
+The medical field demands high accuracy, efficiency, and scalability, traits that align well with the strengths of intelligent agents.
+In this setting, LLM-based agents are beginning to play crucial roles beyond serving as assistive tools Li et al.
+(2024a); Kim et al.
+(2024); Li et al.
+(2025).
 They are involved in essential processes such as diagnosis, medical data analysis, and treatment recommendation.
 By reducing the burden on healthcare professionals and improving clinical outcomes, these agents bring significant value to modern healthcare systems.
-Among various agent system designs, multi-agent systems offer a natural fit for complex medical workflows.
+Among various multi-agent system designs Chen et al.
+(2024c); Liang et al.
+(2024); Wang et al.
+(2024a), they offer a natural fit for complex medical workflows Li et al.
+(2024a); Kim et al.
+(2024).
 Medical imaging and diagnosis tasks generally involve multiple interdisciplinary departments, making coordination and collaboration essential.
@@
 They also help bridge communication gaps between departments and enable standardized procedures across institutions.
-However, current multi-agent systems face a key limitation in workflow design, relying on static, hand-crafted architectures tailored by domain experts.
-While effective in limited scenarios, these fixed workflows lack adaptability to evolving diagnostic techniques or new imaging modalities, requiring time-consuming redesigns.
+However, current multi-agent systems are constrained by how their workflows are designed.
+Existing systems typically rely on static architectures that are hand-crafted Li et al.
+(2024a); Kim et al.
+(2024) by domain experts based on prior experience.
+Though effective in restricted circumstances, these fixed workflows often fail to adapt to complex real-world diagnostic scenarios.
+For example, when diagnostic techniques advance or new imaging modalities are introduced, it can take considerable time and effort to redesign existing workflows.
 Moreover, for each new task or application scenario, developers are required to create customized workflows from scratch, limiting scalability and slowing down deployment.
-To address the aforementioned challenges, we draw inspiration from the field of Automated Machine Learning (AutoML), particularly neural architecture search (NAS), to explore automated architecture design for medical agents.
+To address the aforementioned challenges, we draw inspiration from the field of Automated Machine Learning (AutoML) Elshawi et al.
+(2019); Zhang et al.
+(2021); Chen et al.
+(2024a), particularly neural architecture search (NAS) Wistuba et al.
+(2019), to explore automated architecture design for medical agents.
 NAS has demonstrated its effectiveness in autonomously discovering high-performing neural architectures across a variety of tasks.
-Analogously, we argue that the automatic design of agent architectures can enhance the adaptability of multi-agent systems in dynamic clinical environments.
-To this end, we conceptualize the agent system as a dynamic, graph-based workflow that evolves based on feedback from large language models (LLMs).
-To facilitate effective workflow evolution, we define a comprehensive search space comprising all valid evolution operations.
-These operations are categorized into three levels: node-level, structural-level, and framework-level.
+Analogously, we argue that the automated design of agent architectures can enhance the adaptability of multi-agent systems in dynamic clinical environments.
+To this end, we represent the agent system as a dynamic, graph-based workflow Zhang et al.
+(2025b); Chen et al.
+(2024b) that evolves based on feedback from LLMs.
+To facilitate effective workflow evolution, we define a comprehensive search space comprising valid evolution operations.
+The operations are categorized into three levels: node-level, structural-level, and framework-level.
 Node-level operations include adding or removing nodes and modifying attributes within individual nodes.
 To enable more sophisticated reasoning, structural-level operations introduce mechanisms such as conditional decision-making, iterative loops, and parallel execution, thereby enriching the logical workflow and enhancing diagnostic capabilities.
-At a higher abstraction level, framework-level operations incorporate advanced paradigms such as Chain of Thought Wei et al.
+At a higher abstraction level, framework-level operations incorporate established agent frameworks such as Chain of Thought Wei et al.
 (2022), Reflexion Shinn et al.
-(2023), and Roundtable Discussion Chen et al.
-(2024b), further improving the system’s reasoning and coordination abilities.
-After defining evolution operations, we first established a naive feedforward workflow with only a single node to interact with the LLM and obtain the diagnostic results.
-As shown in Fig.
-1, when diagnostic errors occur, the agent system can identify the root causes and provide both structural and prompt suggestions for workflow evolution.
+(2023a), and Round Table Chen et al.
+(2024c) to guide the search process, further improving the system’s reasoning and coordination abilities.
+After defining the evolution operations, we first establish an initial baseline workflow with only a single node to interact with the LLM and obtain the diagnostic results.
+As shown in Figure 1, when diagnostic errors occur, the agent system can identify the root causes and provide both structural and prompt suggestions for workflow evolution.
 As more cases are processed, the workflow undergoes iterative refinement, progressively incorporating new nodes, structural logic, and framework-level reasoning strategies.
-Over time, it leads to the emergence of a more accurate, robust, and efficient diagnostic system.
+Over time, this results in a more accurate, robust, and efficient diagnostic system.
 This continuous evolution strategy equips the agent system with the capacity for self-improvement to handle ever-changing real-world complexities.
-To validate the effectiveness of our approach, we conduct comprehensive experiments on two medical diagnostic benchmarks.
-The results demonstrate that our method can automatically generate more efficient and accurate agent architectures, significantly outperforming strong baselines.
+To validate the effectiveness of our approach, we conduct comprehensive experiments on two medical diagnostic benchmarks Naqvi (2023); Daneshjou et al.
+(2023).
+The results demonstrate that our method can automatically generate more efficient and accurate agent architectures, significantly outperforming baselines.
 Our main contributions are as follows:
@@
 2 Related Work
-2.1 LLM-based Agent in Medical Field
-In recent years, the reasoning capabilities of large language models (LLMs) have improved significantly Chen et al.
-(2025), leading to their increasing integration into the medical domain to enhance both diagnostic accuracy and treatment effectiveness Hager et al.
+2.1 LLM-Based Agents in the Medical Domain
+In recent years, the reasoning capabilities of LLMs have improved significantly Chen et al.
+(2025), enabling their broader application in the medical domain to enhance diagnostic accuracy and treatment effectiveness Hager et al.
 (2024); Schmidgall et al.
@@
 Meanwhile, MMedAgent Li et al.
-(2024) integrates toolchains to enable multi-modal task execution, supporting robust performance across varied medical applications.
-Additionally, some workKim et al.
-(2024); Li et al.
-(2023) focus on multi-agent collaboration improving efficiency in complex medical workflows.
-Nevertheless, despite their strengths in specific task performance, current systems pay limited attention to the dynamics workflows, which are essential for adapting to the complexity and variability of real-world medical environments.
-2.2 Agent Collaboration and Self-Improving
-Recent studies emphasize the growing importance of agent collaboration and self-improvement in multi-agent systems (MAS) based on large language models (LLMs).
+(2024a) integrates toolchains to enable multi-modal task execution, supporting robust performance across varied medical applications.
+More recent efforts such as MDAgents Kim et al.
+(2024) explore multi-agent collaboration to improve efficiency in complex medical workflows.
+Despite their strengths in specific task performance, current systems pay limited attention to the dynamic workflows, which are essential for adapting to the complexity and variability of real-world medical environments.
+2.2 Multi-Agent Collaboration and Evolution
+Recent studies emphasize the growing importance of agent collaboration and self-improvement in multi-agent systems based on LLMs.
 Common architectures include peer-to-peer setups, centralized controllers, and hierarchical frameworks Chen et al.
@@
 (2025).
-Collaborative and self-improving MAS have shown promise in diverse application domains, like MetaGPTHong et al.
+Collaborative and self-improving multi-agent systems have shown promise in diverse application domains such as MetaGPT Hong et al.
 (2024), LLM-Blender Jiang et al.
@@
 Beyond these domains, recent work focuses on enhancing adaptability in open-ended settings by exploring self-evolving mechanisms: EvoMAC Hu et al.
-(2024) introduces a self-evolving collaboration network trained via textual backpropagation for software-level development tasks.
+(2024) introduces a self-evolving collaboration network that mimics the backpropagation mechanism in neural networks.
 AutoAgents Chen et al.
-(2024a) supports automatic agent generation and flexible collaboration.
+(2024b) supports automatic agent generation and flexible collaboration.
 AFLOW Zhang et al.
-(2025) uses graph search to optimize task execution paths.
+(2025b) uses graph search to optimize task execution paths.
 3 Methodology
 Figure 2: Illustration of node attributes and hierarchical search space.
-(a) Demonstrates the attributes of a basic node and how it can evolve into complex workflow structures such as conditional branches, loops, and parallel executions.
-Each basic node contains configurable parameters including system prompts, human prompts, descriptions, and optional image paths.
+(a) Demonstrates the attributes of a basic node and how it can evolve into complex workflow structures.
+Each basic node contains configurable parameters.
 These attributes allow the agent to execute specific medical diagnostic tasks within the workflow.
 (b) Illustrates the proposed hierarchical search space for dynamic workflow design, consisting of node-level, structural-level, and framework-level operations.
-These enable fine-grained node manipulation, control flow structures (e.g., conditionals, loops, parallelism), and emergent high-level reasoning patterns such as Chain-of-Thought, Reflexion, and Round Table Discussion.
-We propose a dynamic framework that leverages large language models to optimize medical workflows.
-In our approach, workflows are represented as graph-based structures comprising heterogeneous node types, each characterized by specific attributes and functional roles.
-We first introduce the graph-based representation of medical workflows, where each node encapsulates a distinct operation with semantic and procedural properties.
-We then detail how combinations of these nodes define a structured and constrained search space, effectively capturing the spectrum of valid workflow configurations.
-Finally, we present the Workflow Evolution process to enabling continuous improvement in workflow design and performance.
+These enable fine-grained node manipulation, control flow structures, and emergent high-level reasoning patterns.
+We propose a dynamic framework that leverages LLMs to optimize medical agent workflows.
+In our approach, workflows are represented as graph-based structures Zhang et al.
+(2025b); Chen et al.
+(2024b), designed to include two node types: tool nodes and basic nodes.
+We first introduce the functions and attributes of different node types.
+We then formulate a structured and constrained search space, which defines the set of structural operations available for workflow evolution.
+Finally, we describe the workflow evolution process, which analyzes workflow shortcomings based on diagnostic errors and generates actionable suggestions for improvement.
 3.1 Workflow Node Definition
-In this section, we define the concept of a medical workflow and its foundational components, which collectively form the basis for dynamic task execution within the proposed framework.
-A workflow is modeled as a sequence of interconnected nodes, categorized into basic and tool types.
-A basic node directly interacts with the LLM, generating a single result based on the provided prompts.
-In contrast, a tool node utilizes predefined tools accessible to the LLM, with an emphasis on supporting workflow evolution.
-For simplicity, only one tool, the image search tool, is introduced.
-This tool accepts an image as input and returns the top k most similar disease labels along with their similarity scores, which serve as diagnostic references for the LLM.
-Basic nodes serve as the foundation from which other structures evolve; for example, Figure  2 (a) show that a basic node combined with a looping condition transforms into a loop node, while one with a conditional branch becomes a conditional node.
-Multiple basic nodes running concurrently create a parallel structure.
-Each structural type plays a distinct role in facilitating the flexible and efficient execution of complex medical workflows.
-For the basic node type, the following key attributes are included:
+Our medical agent workflow is formalized as a graph Zhang et al.
+(2025b); Chen et al.
+(2024b), where nodes represent specialized agents responsible for distinct functional roles, and edges encode the execution order between them.
+Nodes are categorized into two types: basic nodes and tool nodes.
+We describe each node type in detail below.
+Basic node.
+A basic node directly interacts with the LLM by processing system and human prompts to generate task-specific responses.
+It serves as the fundamental building block of medical agent workflows.
+As illustrated in Figure 2(a), each basic node is configured with the following key attributes:
+  • Node Name: The name of the basic node.
   • System Prompt: Defines the role or identity of the agent within the node.
-  • Human Prompt: Specifies the instructions for the agent to carry out the designated task.
-  • Description: Provides a concise explanation of the node’s function or purpose.
-For our tool node type, the following key attributes are included:
+  • Human Prompt: It specifies the instructions for the agent to carry out the designated task.
+  • Description: It provides a concise explanation of the node’s function or purpose.
+Tool node.
+A tool node can invoke predefined external tools.
+For example, image searching retrieves supplementary information that enhances the agent’s reasoning and decision-making.
+The tool node includes the following key attributes:
+  • Node Name: The name of the tool node.
   • Tool Name: The name of the tool the node will use.
-  • Description: Provides a concise explanation of the node’s function or purpose.
+  • Description: It provides a concise explanation of the node’s function or purpose.
+As a concrete example, we define an image search tool.
+This tool retrieves similar medical images from a reference database by leveraging the pre-trained CLIP (ViT-L/14) Radford et al.
+(2021) model to extract and match image feature vectors without fine-tuning.
+Given a query image, it returns the top-k most similar cases with associated disease labels, which serve as diagnostic references for the LLM.
 3.2 Search Space
-To facilitate systematic workflow evolution, we propose a hierarchical and expressive search space that defines actionable strategies for agents across node-level, structural-level, and framework-level operations, enabling precise adjustments to workflow architecture and task-specific parameters.
-Node-Level Operations. Node-level operations focus on modifying individual workflow components to optimize functionality and streamline execution.
-  • Add: Introduces a new node to the workflow, which can be a basic functional node or a tool node integrated with specialized capabilities.
-By adding nodes, workflows can accommodate additional capabilities, enhancing problem-solving capabilities.
-  • Remove: Eliminates unnecessary or redundant nodes to improve workflow efficiency. Removing nodes helps streamline execution paths, resulting in a more concise and targeted process.
-  • Modify: Adjusts prompt configurations for existing nodes. This enables fine-grained control over node behavior to better fulfill task requirements.
-Structural-level Operations. Structural operations enable the modification and expansion of overarching workflow logic, addressing branching, feedback loops, and parallel execution paths.
-  • Conditional Structures: Conditional branches redirect workflow execution based on specified criteria, creating dynamic paths for tailored task handling.
-This adds flexibility and adaptability to workflows by allowing different paths based on decision logic and runtime conditions.
-  • Loop Structures: Feedback loops allow iterative node execution under specific conditions, supporting tasks that require repeated evaluation or refinement.
-Proper loop designs optimize iterative processes, including condition thresholds, iteration limits, and feedback paths.
-  • Parallel Structures: Parallel execution paths enable concurrent multi-agent analysis, enhancing computational efficiency and fostering collaborative workflows.
-These structures are particularly beneficial when tasks can be divided into independent sub-tasks that can be processed simultaneously, thereby reducing the overall execution time.
-Situations that demand high throughput, involve large data sets, or require simultaneous execution of independent processes can greatly benefit from parallel structures.
-Additionally, parallel execution allows different agents to express their viewpoints independently without being influenced by others, promoting unbiased contributions.
-Once all agents have shared their perspectives, they can proceed to integrate and consider each other’s insights, leading to a more comprehensive and collaborative outcome.
+To enable the discovery of better workflows, we propose a hierarchical search space that defines actionable strategies for agents across node-level, structural-level, and framework-level operations.
+As illustrated in Figure 2(b), this design allows for precise adjustments to both workflow structure and node prompts.
+Node-Level Operations.
+Node-level operations involve modifying individual nodes to enhance both the functionality and efficiency of the workflow.
+Specifically, we define the following three types of operations:
+  • Add: This operation adds a new node to the workflow, which can be either a basic node or a tool node.
+By adding such nodes, workflows can accommodate new functionalities and improve their reasoning and diagnostic performance when the complexity of tasks increases.
+  • Remove: It eliminates unnecessary or redundant nodes to improve workflow efficiency. Removing nodes helps streamline execution paths, resulting in a more concise and targeted process.
+  • Modify: This operation can update the prompt configurations of existing nodes, making the prompts more precise and better aligned with task requirements.
+Structural-Level Operations.
+Structural operations enable modifications to the workflow’s execution structure, including conditional branches, iterative loops, and parallel execution paths.
+These modifications enable the medical agent workflow to evolve from a simple linear sequence into a more adaptive and structured framework that more effectively supports complex diagnostic tasks.
+We define three types of structural operations.
+  • Conditional Structures: Conditional structures allow the workflow to take different execution paths based on logical conditions.
+This supports adaptive behavior by allowing the agent to select appropriate actions according to context, such as the patient symptoms or diagnostic outcomes.
+  • Loop Structures: Loop structures enable repeated execution of a node or sub-workflow until a termination condition is satisfied, such as achieving stable outputs or reaching a maximum iteration count.
+Additionally, loop structures support error correction and progressive refinement by allowing the medical workflow to revisit and improve prior decisions based on feedback.
+  • Parallel Structures: Parallel structures enable multiple nodes to conduct independent analyses of the same case simultaneously, each focusing on different clinical aspects such as image symptom description and differential diagnoses.
+By integrating these diverse perspectives through consensus or fusion mechanisms, the workflow achieves a more comprehensive and reliable diagnostic outcome.
+Figure 3: Workflow evolution over iterations.
+We visualize the diagnostic workflow at three stages: initial (Iteration 0), intermediate (Iteration 3), and final (Iteration 6).
+The evolution shows incremental improvements, including adding conditional branches, parallel paths, enhancing diagnostic accuracy and robustness through iterative refinement.
 Framework-Level Design.
-The search space implicitly includes high-level conceptual workflow patterns—such as roundtable discussions—as abstract templates.
-These patterns are not pre-defined as fixed modules, but rather emerge dynamically through the composition of lower-level nodes and structural operations during the search process, as shown in figure2.
-For instance, the behavior of a roundtable discussion can be simulated by orchestrating a set of interconnected nodes in a looped structure.
-This design allows the framework to flexibly construct complex task-specific workflows by synthesizing reusable primitives in a goal-driven manner.
-  • Reflection Framework: This framework focuses on iterative self-assessment and knowledge refinement to boost problem-solving abilities through reflective thinking.
-Promoting continuous evaluation and learning allows agents to adapt and improve over time, leading to more informed decisions.
-  • Collaboration Framework: By allowing experts from various fields to work concurrently, this framework enhances the ability to address complex problems with better solutions.
-  • Chain-of-Thought Reasoning: It decomposes complex tasks into sequential, modular steps, allowing for detailed analysis at each stage.
-By systematically addressing each component, it enhances the accuracy and thoroughness of the overall execution process.
-  • Round Table Discussion Framework: Multiple agents collaborate and deliberate around a common task or theme to achieve a more efficient or accurate solution.
-Each agent may employ distinct models or algorithms and share information with one another, akin to a round table meeting, to integrate diverse insights and optimize the final outcome.
-  • Conquer-and-Merge Discussion (CMD) Framework: Parallel analytical groups with voting mechanisms for consensus driven decision making.
+At the framework-level, the search space incorporates existing collaborative frameworks methods, including Chain of Thought Wei et al.
+(2022), Reflexion Shinn et al.
+(2023a), Round Table Chen et al.
+(2024c), and Conquer-and-Merge Discussion (CMD) Framework Wang et al.
+(2024a).
+Rather than directly adopting these structures, our approach dynamically composes customized collaboration patterns through combinations of basic nodes and structural operations.
+To illustrate this process, we take the Round Table framework Chen et al.
+(2024c) as an example.
+This framework organizes multiple expert nodes into a structured discussion process, where each node represents a specialized role (e.g., dermatologist, pathologist).
+In the first round, all expert nodes generate initial diagnostic suggestions.
+In subsequent rounds, the experts take turns refining their opinions in a sequential manner, using the previous responses and reasoning from other experts as context.
+This enables our framework to dynamically construct workflow structures that are functionally similar to established multi-agent frameworks, while allowing these structures to be customized at both the node and structural levels to meet the specific requirements of the current diagnostic task.
 3.3 Workflow Evolution
-Our proposed workflow evolution process is a continuous iterative cycle.
-Initially, large language models perform a diagnostic analysis of the workflow execution to identify potential issues, categorizing them into various error cases.
+Our proposed workflow evolution process is an iterative refinement loop in which the LLM analyzes diagnostic failures from previous executions, identifies their root causes, and classifies them into two distinct error types: Image Understanding Errors and Diagnostic Errors.
+This classification guides targeted improvements to both the workflow structure and node prompts.
 The causes of these errors are collected and analyzed to generate improvement suggestions.
-These suggestions are then transformed into actionable workflow modifications, aligning with the requirements to ensure targeted enhancements.
-The updated workflow is validated and subsequently updated, after which the iteration continues with the next diagnostic phase.
-This process of diagnosing, collecting causes, analyzing, receiving suggestions, verifying, and updating ensures that the workflow is continually refined and optimized for both structure and execution performance.
+These suggestions are then transformed into actionable workflow modifications.
+The modified workflow is validated and, if successful, deployed for the next diagnostic iteration.
+Through iterative cycles of diagnosis, analysis, feedback, and refinement, the workflow is continually improved and optimized.
+To better understand this evolution process, we provide a detailed description of the entire mechanism in the following.
 First, LLM focuses on diagnosing workflow errors, and it aggregates outputs from all workflow nodes and classifies errors into two primary categories.
-Image Understanding Errors: Issues arising from inaccuracies in feature recognition or extraction, which compromise the accuracy of diagnostic decisions.
-Diagnostic Decision Errors: Mistakes emerging from flawed reasoning or incorrect decision-making, even when image features are correctly identified.
-Then, LLM traces the root causes of these errors, determining whether they originate from inaccurate image feature recognition or deficiencies in diagnostic reasoning.
-Critical data related to the problematic nodes is preserved during this process, enabling a more in-depth analysis to uncover the origins of diagnostic inaccuracies.
-Based on this analysis, the system proceeds to analyze the workflow’s logical structure and extract detailed descriptive metadata from its components.
-This process is facilitated by the visual perception capabilities of the large language model, which enable the generation of a structured, graph-based representation of the workflow.
-The resulting representation provides both a visual and logical overview of the workflow architecture, clearly illustrating the interconnections and dependencies among components.
-The workflow is visualized using Mermaid, which allows for direct viewing of loop conditions and conditional branches within the generated flowchart.
-This structural representation allows the system to accurately identify the relationships between nodes and understand the specific functional roles they fulfill within the overall process.
-By integrating this structural understanding with an analysis of key issues, the LLM can offer more targeted and effective suggestions for improving the workflow.
-Next these suggestions undergo a filtering process to eliminate impractical proposals, such as those requiring external inputs beyond the scope of the intelligent agent’s capabilities.
-The remaining suggestions are further categorized into two groups: Structural suggestions: Suggestions aimed at optimizing the overall workflow architecture and structural dependencies.
-Prompt suggestions: Suggestions focused on refining task-specific configurations, such as prompt templates, to enhance node performance.
-Finally, these filtered suggestions are refined and integrated into the workflow to support effective evolution and optimization.
-Initially, the suggestions are reformulated into actionable updates and returned in standardized formats as JSON.
-The structural suggestions are primarily aimed at improving the overall architecture of the workflow.
-Due to their potential impact on the underlying structure, these suggestions must undergo a rigorous validation process.
-This step is essential to ensure that no invalid or inappropriate actions from the search space are introduced, as such actions could disrupt or compromise the integrity of the existing workflow design.
-Once validated, they are applied to optimize the workflow’s architecture and execution flow.
-Prompt suggestions, on the other hand, directly modify task-specific configurations, such as prompt templates, and are applied seamlessly without requiring additional validation steps.
-By systematically refining and integrating these suggestions, the workflow evolves iteratively, achieving continuous improvement in both its structure and diagnostic performance.
-Through validation and integration, the framework systematically incorporates improvements while preventing the introduction of inefficient or erroneous actions that could disrupt the workflow structure.
-This refinement mechanism ensures that every validated structural recommendation and prompt modification contributes to meaningful enhancements in workflow functionality, enabling continuous and targeted optimization.
-The entire process—from diagnosis and error analysis to workflow updates—can be viewed as an evolutionary cycle.
-During each iteration, we run multiple examples and then refine the workflow to produce an updated version.
-This process is repeated continuously until the final workflow’s accuracy on the validation set converges, signifying the completion of the evolution.
+Image Understanding Errors: Issues arising from inaccuracies in image descriptions, which compromise the accuracy of diagnostic decisions.
+Diagnostic Errors: Mistakes emerging from flawed reasoning or incorrect decision-making, even when the medical image descriptions are correct.
+Then, LLM traces the root causes of these errors, determining whether they originate from inaccurate image descriptions or errors in diagnostic inference.
+Next, we need to analyze the structure of the workflow.
+Textual descriptions are commonly used to record workflow structures.
+However, they often fail to clearly represent the workflow’s execution logic in a way that is easily interpretable by LLMs.
+In particular, the recorded node and edge relationships are not organized according to the workflow’s execution sequence, but instead appended incrementally as structural updates are made.
+This can make it difficult for LLMs to accurately interpret the workflow structure, potentially leading to incorrect analysis of the execution logic.
+To address this limitation, we convert the recorded structural information into Mermaid Knut Sveidqvist (2024) code and generate a corresponding visual workflow diagram.
+This visualization provides a clearer and more structured view of the workflow, enabling the LLM to better understand its execution logic and offer more accurate suggestions for improvement.
+Having analyzed the workflow structure and the issues identified by the LLM, we now explain how the LLM translates these insights into actionable modifications for iterative refinement.
+Our framework implements a systematic refinement process to ensure safe and effective workflow evolution.
+First, generated suggestions are filtered to discard unrealistic actions, such as those that require external inputs beyond the agent’s capabilities, retaining only feasible and actionable ones.
+These filtered suggestions are then categorized into two types: Structural suggestions, which aim to optimize the workflow’s architecture (e.g., adding loops or parallel branches), and Prompt suggestions, which refine task-specific configurations such as node prompts to improve reasoning quality.
+To enable automated processing, all suggestions are reformulated into standardized formats that clearly specify the intended modifications.
+Before processing, we define a set of prompt templates that outline the required parameters for the different types of node modifications.
+Based on these templates, the LLM reformulates each suggestion into a structured format, which provides the structured data required for updating the workflow.
+Structural suggestions are carefully validated before integration, as they may break the normal execution of the workflow.
+To prevent such modifications, the framework incorporates a set of structural validation mechanisms to ensure that all modifications maintain the integrity of the workflow graph and its execution logic.
+To prevent duplicate edges, the workflow rejects any modification that attempts to add a new edge between two nodes that are already connected.
+This prevents redundant or conflicting execution paths.
+Additionally, the framework also limits the maximum number of iterations to prevent infinite loops and validates the directionality of edges when introducing cycles, ensuring that loops have a well-defined exit condition.
+These validation mechanisms work together to safeguard the workflow from invalid or disruptive changes, ensuring that all structural modifications are both logical and executable.
+Prompt suggestions only affect the internal configuration of individual nodes and do not alter the workflow structure.
+As a result, they can be applied directly without additional validation.
+These modifications typically include updates to the system prompt, which provides a more accurate definition of the node’s role or identity, and the human prompt, which clarifies the specific task a node should complete.
+The entire process, from diagnosis and error analysis to workflow updates, can be viewed as an evolutionary cycle.
+During each iteration, multiple disease cases are processed, and the workflow is subsequently refined to produce an updated version.
+This cycle is repeated iteratively until the accuracy of the final workflow on the validation set converges, indicating the completion of the evolution process.
 4 Experiments
 4.1 Experimental Setup
-Figure 3: Workflow evolution over iterations.
-We visualize the diagnostic workflow at three key stages: the initial state (Iteration 0), an intermediate state after partial refinement (Iteration 3), and the final evolved workflow (Iteration 6).
-The evolution process demonstrates how the system incrementally improves its structure by adding conditional branches, parallel reasoning paths, and error-specific loops.
-These structural refinements enable more accurate and robust diagnostic reasoning through iterative self-adaptation.
 Datasets.
-To evaluate the effectiveness of our proposed framework, we select the task of skin disease diagnostics as the primary testbed.
-This domain is particularly suitable for evaluating multi-agent architectures due to its inherent complexity: accurate diagnosis often requires interdisciplinary knowledge and a multi-step reasoning process involving visual inspection of skin lesions, patient history, and symptom-diagnosis correlations.
-We utilized two dermatological datasets to evaluate the workflow evolution:
-SKIN Concepts Dataset.
-SKINCONDaneshjou et al.
-(2023) builds upon the Fitzpatrick 17k skin disease datasetGroh et al.
-(2022, 2021), including 3230 images annotated with 48 clinically relevant concepts, such as "plaque," "scale," and "erosion." These annotations span 113 types of skin diseases and are based on established clinical vocabulary used by dermatologists, ensuring their applicability across diverse disease processes.
-Augmented Skin Conditions Image Dataset.
-Skin Conditions Naqvi (2023)contains a total of 2,394 augmented images across six different dermatological conditions, with each category comprising 399 images.
-This balanced structure makes it ideal for training machine learning models, particularly in the field of medical image analysis.
-Data Processing.
-During preliminary data preparation, we identified a subset of images within the datasets that depicted skin conditions located in private or sensitive body areas.
-Due to content moderation constraints, these images were rejected by the LLM and could not be processed during workflow execution.
-Additionally, we encountered another issue where approximately 20 images had broken URL links, preventing them from being downloaded.
-To maintain the integrity of our experiments and ensure consistent, reliable evaluation across all agent architectures, we excluded these samples from the dataset.
-Datasets Split: SKIN Concepts Dataset: The training and validation sets contain 50 randomly sampled images each, and the test set includes 224 images, ensuring that all disease types are represented with two images per class.
-SKIN Conditions Dataset: Training and validation sets each have 50 images; the test set contains 120 images, with 20 per condition.
-Implementation Details. We implemented an experimental process that incorporates training, validation, and testing phases.
-After each training batch, the evolutionary mechanism adjusts the workflow structure.
-The model’s performance is then validated on a separate set.
-This cycle of training and evolution repeats until training completion, followed by a final evaluation of the test set.
-Our framework is built on LangGraphInc.
+To evaluate the effectiveness of our proposed framework, we conduct experiments on two dermatological image classification datasets Daneshjou et al.
+(2023); Naqvi (2023) to assess workflow evolution.
+Dermatology presents a challenging and representative domain for evaluating multi-agent architectures, as accurate diagnosis typically demands both specialized visual analysis and integrative reasoning.
+Below is a detailed description of the two datasets.
+Skin Concepts Daneshjou et al.
+(2023) extends the Fitzpatrick 17k skin disease dataset Groh et al.
+(2022, 2021) and includes 3,230 images annotated with 48 clinically relevant skin concepts.
+For evaluation, we adopt a split of 50 images for training, 50 for validation, and 224 for testing, ensuring two test images per disease type to maintain balanced coverage across all categories.
+Augmented Skin Conditions Naqvi (2023) consists of 2,394 synthetically augmented images across six dermatological conditions, with 399 images per class.
+For evaluation, we adopt a similar split of 50 training, 50 validation, and 120 test images, with 20 images per disease type, to ensure sufficient and balanced evaluation of diagnostic performance.
+During data processing, we observed that the LLM rejected a small number of images depicting sensitive body areas due to content moderation policies.
+Additionally, some image data with broken URLs could not be downloaded.
+To ensure experimental integrity and enable fair and consistent evaluation across all agent architectures, we excluded these samples from our experiments.
+In total, 100 images from the Skin Concepts Daneshjou et al.
+(2023) dataset were removed.
+Figure 4: Top-k accuracy improves with workflow iterations, and convergence signals system stabilization using GPT-4o.
+Baselines.
+To establish a comparative framework for evaluating the effectiveness of our workflow evolution methodology, we implemented three baseline approaches.
+The first baseline utilizes LLM to give a direct diagnosis of input images, without any intermediate reasoning or workflow adjustments.
+This provides a straightforward benchmark for assessing diagnostic accuracy.
+The second baseline augments the LLM with a CoTWei et al.
+(2022) prompting mechanism.
+This approach guides LLM to reason through intermediate steps before producing the final diagnosis.
+The third baseline adopts the Round Table Chen et al.
+(2024c) framework, a multi-agent collaboration paradigm inspired by human expert panels.
+In this setup, multiple agents take turns to refine a shared solution through iterative deliberation.
+Each agent reviews the reasoning history and outputs of previous agents, then contributes an updated response based on its own perspective.
+This process continues for a fixed number of iterations or until convergence, aiming to achieve a more comprehensive and accurate final decision.
+Implementation Details.
+Our experimental setup includes training, validation, and testing phases.
+During training, the workflow evolves through iterative refinement based on diagnostic outcomes.
+Each updated workflow is validated after every batch to assess performance improvements.
+Once training completes, the final workflow is evaluated on the test set to measure overall performance.
+Our framework extends LangGraph Inc.
+(2023) with customized node designs to support flexible and modular workflow construction.
+To facilitate extensibility, we introduce a configurable tool list that allows users to easily add or integrate their own custom tools.
+As a built-in example, we include an image search tool designed to find visually similar images.
+We use the pre-trained CLIP (ViT-L/14) Radford et al.
+(2021) model to extract feature vectors from images that are not included in the training, validation, or test sets.
+These vectors are then stored in the Pinecone Pinecone Systems (2023) vector database to support image similarity searches.
+When a new image is input, its feature vector is computed and compared against the stored vectors to enable fast and efficient similarity searches.
+To evaluate the generality and robustness of our framework across diverse LLMs, we conduct experiments using the following models: GPT-4o-2024-05-13 Achiam et al.
+(2023), Claude-3-5-sonnet-20240620 Anthropic.
+(2024), and GPT-4o-mini-2024-07-18 Menick et al.
 (2024).
-LangGraph provides us with some pre-packaged content, and we only need to develop our node designs on top of LangGraph.
-Additionally, we provide a tool list where you can add your custom tools.
-In the tool list, we have defined an image search tool for finding similar images.
-Image search tool utilizes the pre-trained CLIP(ViT-L/14)Radford et al.
-(2021) model to extract feature vectors without any further fine-tuning from images for data that is not part of the training, validation, and testing sets, storing these vectors in the PineconePinecone Systems (2023) database.
-When searching for similar images using the image tool search, it can directly match the feature vectors stored in this database.
-Figure 4: Top-k diagnostic accuracy over workflow evolution iterations.
-This figure shows how the Top-k accuracy improves as the workflow undergoes iterative optimization.
-Each point represents the diagnostic performance after one round of workflow evolution.
-The increasing trend demonstrates that structural refinements and prompt updates effectively enhance diagnostic capability.
-The curve eventually converges, indicating that the system stabilizes once a sufficiently optimized workflow is discovered.
-The results in the figure were obtained using the GPT-4o.
-Baselines.
-To establish a comparative framework for evaluating the effectiveness of our workflow evolution methodology, we implemented two baseline approaches for image-based dermatological diagnosis.
-The first baseline utilizes GPT-4o as a single-agent model to perform direct diagnosis of input images, without any intermediate reasoning or workflow adjustments.
-This provides a straightforward benchmark for assessing diagnostic accuracy.
-The second baseline augments GPT-4o with a Chain-of-Thought (CoT) prompting mechanism.
-This approach guides GPT-4o to reason through intermediate steps before producing the final diagnosis.
-Although CoT enhances interpretability compared to direct diagnosis, it does not leverage adaptive or evolving workflows, which are central to our proposed methodology.
-Figure 5: Per-disease consensus accuracy (cons@64) across different methods.This figure illustrates the prediction stability of each method under six different disease categories, evaluated via majority-vote consensus over 64 samples per class.
-Our method consistently achieves higher cons@64 scores, indicating improved diagnostic robustness and intra-class agreement across cases.
-The third baseline involves a "round table" discussion setup, which is a type of multi-agent system configuration.
-In this approach, multiple agents collaborate and deliberate around a common task or theme to achieve a more efficient or accurate solution.
-Each agent may employ distinct models or algorithms and share information with one another, akin to a round table meeting, to integrate diverse insights and optimize the final outcome.
-These baselines allow for quantifying the benefits of workflow evolution in terms of generalization ability and diagnostic performance.
+The temperature parameter is set to 1.0 to encourage diverse reasoning paths, and the random seed is 42.
+Table 1: Top-k diagnostic accuracy (%) of different methods using GPT-4o, GPT-4o-mini and Claude 3.5 Sonnet on Skin Concepts and Skin Conditions.
 LLM Method Skin Concepts Accuracy (%) Skin Conditions Accuracy (%) Top-1 Top-3 Top-5 Top-1 Top-3 Top-5 GPT-4o IO 20.27 30.63 36.04 50.83 78.33 86.67 CoT Wei et al.
 (2022) 18.47 28.83 33.78 55.83 76.67 82.50 Round Table Chen et al.
-(2024b) 21.17 27.93 32.43 45.83 75.83 80.83 Ours 29.28 40.09 50.45 90.83 95.00 100.00 GPT-4o-mini IO 11.71 20.72 23.87 27.50 69.17 80.83 CoT Wei et al.
+(2024c) 21.17 27.93 32.43 45.83 75.83 80.83 Ours 29.28 40.09 50.45 90.83 95.00 100.00 GPT-4o-mini IO 11.71 20.72 23.87 27.50 69.17 80.83 CoT Wei et al.
 (2022) 6.31 15.32 24.32 22.50 65.00 84.17 Round Table Chen et al.
-(2024b) 10.81 19.82 23.42 25.83 70.00 78.33 Ours 13.51 21.62 24.77 45.83 74.17 85.83
-Table 1: Top-k diagnostic accuracy (%) of different methods using GPT-4o and GPT-4o-mini on skin concepts and skin conditions. Higher values indicate better performance.
+(2024c) 10.81 19.82 23.42 25.83 70.00 78.33 Ours 13.51 21.62 24.77 45.83 74.17 85.83 Claude 3.5 Sonnet IO 17.12 24.77 26.13 40.00 68.33 75.83 CoT Wei et al.
+(2022) 14.86 22.07 24.32 36.67 63.33 73.33 Round Table Chen et al.
+(2024c) 15.77 22.52 26.58 35.00 66.67 74.17 Ours 28.83 35.14 38.29 95.83 98.33 99.17
+Evaluation Metrics.
+To evaluate diagnostic models, we used Top-1, Top-3, and Top-5 accuracy metrics, comparing predictions against ground truth labels.
+During validation, these accuracies were tracked across workflow evolution to observe improvements, while in testing, final accuracies were compared with baselines to highlight the benefits of adaptive workflows.
+In addition to accuracy, we assess the consistency of predictions across repeated inputs.
+We adopt the majority voting consensus metric, denoted as cons@64 Guo et al.
+(2025), which reflects how often a method arrives at a consistent diagnosis across multiple inputs of the same class (n=64).
+This metric offers a proxy for model robustness and reliability under input variance and sampling noise.
+These metrics allow us to quantify both the progression of accuracy during workflow evolution and the overall generalization ability of the proposed method when tested on unseen data, particularly in diverse disease categories.
+4.2 Results
+Evolution Visualization.
+To better understand how our workflow evolves, we visualize the structure at three representative stages: the initial workflow (Iteration 0), an intermediate state (Iteration 3), and the final evolved workflow (Iteration 6).
+Figure 3 presents a comparative view of these stages.
+These visualizations reveal how the system progressively incorporates structural refinements, such as conditional branches, parallel reasoning paths, and error-specific loops.
+Cost and Complexity Analysis.
+The results in Table 2 show that the framework evolves into more complex workflows by the end of training.
+The final workflows used in testing contain more nodes and branches than the training-phase averages, indicating an increase in both structural complexity and token consumption driven by the evolutionary process.
+Table 2: Resource Consumption and Workflow Complexity. This table summarizes resource consumption and workflow complexity across different metrics during training (Evo) and testing phases.
+Metric Train (Evo) Test
+Number of Nodes 8.7 11
+Number of Branches 4.0 5
+Time (s) 210 621
+Total Tokens (k) 361 349
+Performance Over Iterations.
+To quantitatively evaluate the effectiveness of our workflow evolution framework, we track the diagnostic performance, measured by Top-1 accuracy, Top-3 accuracy, and Top-5 accuracy across multiple evolution iterations.
+At each iteration, the system refines its workflow based on prior diagnostic errors and structural improvements, allowing it to incrementally enhance its reasoning capabilities.
+Figure 4 illustrates the iterative enhancement in accuracy as the workflow progresses, compared to the baseline accuracy achieved by directly inputting images into the LLM for diagnosis (IO line).
+The evolved workflow consistently surpasses the baseline performance, exhibiting an initial upward trend attributed to structural refinement and prompt optimization.
+This is followed by convergence as the system stabilizes into an effective configuration.
+This pattern underscores the framework’s capacity to learn from experience and self-optimize over time.
+In the context of the Skin Concepts dataset Daneshjou et al.
+(2023), the evolved workflow demonstrates robust generalization capabilities, achieving commendable results across all categories in the testing set.
+Evolution Results Analysis.
+To comprehensively evaluate the effectiveness of our method in diagnostic tasks, we compare it against various baselines as shown in Table 1.
+Specifically, we report the Top-k diagnostic accuracy on dermatological tasks using the Skin Concepts Daneshjou et al.
+(2023) and Skin Conditions Naqvi (2023) datasets across three different LLMs.
+Our approach consistently outperforms existing baselines, including IO, CoT Wei et al.
+(2022), and Round Table Chen et al.
+(2024c), across all LLMs and evaluation metrics.
+Our method achieves superior results in Top-1, Top-3, and Top-5 accuracies for both datasets.
+These results highlight the adaptability and robustness of our method across different LLMs and diverse medical diagnostic scenarios.
+The consistent superiority in Top-k accuracies underscores the effectiveness and reliability of our evolved workflow, making it a promising solution for improving diagnostic performance.
+Figure 5: Per-disease consensus accuracy (cons@64) across different methods.
+This figure shows prediction stability across six disease categories, evaluated via majority-vote consensus over 64 samples per class.
+Our method achieves higher cons@64 scores, indicating greater diagnostic robustness and consistency.
+Workflow Stability.
+In addition to diagnostic accuracy, we assess the stability of our evolved workflows by evaluating their consensus behavior across multiple independent samples.
+We report per-class consensus accuracy (cons@64 Guo et al.
+(2025)) for each method using the GPT-4o model on the Skin Concepts dataset Daneshjou et al.
+(2023), as shown in Figure 5.
+Our method consistently achieves higher consensus scores across a wide range of conditions.
+4.3 Ablation Analysis
+We conduct a comprehensive ablation study to evaluate the contribution of individual workflow modification actions in our evolution framework.
+Actions are classified into three categories: Add, Modify, and Remove.
+In the study, we assess the impact of three key operations: adding tool nodes, modifying node prompts, and removing nodes.
+By excluding each operation from the search space, we observe its contribution to overall performance.
+Table 3 shows that adding tool nodes and modifying prompts notably improve accuracy, while removing nodes has minimal impact on accuracy but enhances execution efficiency by reducing redundancy.
+Table 3: Ablation results showing Top-k accuracies when specific operations are disabled. Parentheses indicate accuracy drops relative to the full model.
 Operation Top-1 Acc.
@@
 (%) Add Tool Node 21.62 (-7.66) 30.18 (-9.91) 36.94 (-13.51) Modify Node Prompt 19.37 (-9.91) 27.93 (-12.16) 33.78 (-16.67) Remove Node 28.83 (-0.45) 41.44 (+1.35) 50.90 (-0.45)
-Table 2: Ablation results of different operations.
-Each row shows the Top-1, Top-3, and Top-5 accuracies when a specific operation is disabled.
-Values in parentheses indicate accuracy drop compared to the full model Top-1(29.28.%), Top-3(40.09%), and Top-5(50.45%).
-Evaluation Metrics.
-To evaluate the diagnostic models, we adopted accuracy-based metrics across Top-1, Top-3, and Top-5 predictions.
-Each diagnostic attempt produces a ranked set of predictions based on the model’s confidence scores, and the correctness of these predictions is compared against ground truth labels.
-During the validation phase, the Top-1, Top-3, and Top-5 accuracies are measured after each evolution iteration, enabling us to observe how these metrics improve as the workflow evolves.
-In the test phase, the final accuracies obtained with our workflow evolution method are compared against those of the baselines, demonstrating the advantages of dynamic and adaptive workflows.
-Following Wang et al.
-(2023), we adopt the majority voting consensus metric, denoted as cons@64, which reflects how often a method arrives at a consistent diagnosis across multiple inputs of the same class (n=64).
-This metric offers a proxy for model robustness and reliability under input variance and sampling noise.
-These metrics allow us to quantify both the progression of accuracy during workflow evolution and the overall generalization ability of the proposed method when tested on unseen data, particularly in diverse disease categories.
-4.2 Analysis of Evolutionary Processes
-Evolution Visualization.
-If the workflow nodes are directly input to the LLM in text form, the LLM may easily overlook structural relationships such as loops and conditional branches.
-To better understand how our workflow evolves over time, we visualize the structure at three representative stages: the initial workflow (Iteration 0), an intermediate state (Iteration 3), and the final evolved workflow (Iteration 6).
-These visualizations reveal how the system progressively incorporates structural refinements, such as conditional branches, parallel reasoning paths, and error-specific loops.
-Figure 3 presents a comparative view of these stages.
-Quantitative Analysis To quantitatively evaluate the effectiveness of our workflow evolution framework, we track the diagnostic performance—measured by Top-1 accuracy, Top-3 accuracy and Top-5 accuracy—across multiple evolution iterations.
-At each iteration, the system refines its workflow based on prior diagnostic errors and structural improvements, allowing it to incrementally enhance its reasoning capabilities.
-Figure 4 illustrates how the Top-1 accuracy, Top-3 accuracy, and Top-5 accuracy change as the workflow evolves.
-The straight line represents the accuracy of directly inputting images into the LLM for diagnosis.
-It can be observed that the accuracy of the evolved workflow is always higher than the IO line.
-We observe a consistent upward trend during the early iterations, indicating that the refinement of workflow structure and prompt configurations leads to improved diagnostic performance.
-The accuracy eventually converges, suggesting that the system has reached a stable and effective workflow configuration.
-This pattern demonstrates the framework’s ability to learn from experience and self-optimize over time.
-Additionally, in the Skin Concepts dataset, both the training and validation sets consist of 50 randomly selected images, while the testing set includes images from all categories.
-Our workflow iteratively evolves on the training set and is validated on the validation set.
-Ultimately, it achieves good results on the testing set, which includes all categories.
-This demonstrates the generalization capability of our evolved workflow.
-4.3 Performance Comparison Results
-Diagnostic Accuracy Comparison.
-To comprehensively evaluate the effectiveness of our method in dermatological diagnostic tasks, we compare various prompting strategies across multiple state-of-the-art large language models (LLMs).
-Specifically, we report the Top-k diagnostic accuracy on both skin concepts and skin conditions using GPT-4o and GPT-4o-mini.
-As shown in Table 1, our approach consistently outperforms existing baselines, including IO, CoT Wei et al.
-(2022), and Round TableChen et al.
-(2024b), achieving superior results across all LLMs and evaluation metrics.
-As shown in Figure 1, our method achieves substantial gains in Top-k accuracy, highlighting its superior diagnostic capability and robustness across different models.
-Workflow Stability.
-In addition to diagnostic accuracy and structural adaptability, we assess the stability of our evolved workflows by evaluating their consensus behavior across multiple independent samples.
-We report per-class consensus accuracy (cons@64) for each method using the GPT-4o model on the Skin Concepts dataset, as shown in Figure 5.
-Our method consistently achieves higher consensus scores across a wide range of conditions, including both common (e.g., eczema, psoriasis) and rare or visually complex diseases (e.g., impetigo, rosacea).
-This demonstrates not only the general robustness but also the fine-grained reliability of the evolved workflows.
-4.4 Component Ablation Analysis
-We conduct a comprehensive ablation study to investigate the role of different workflow modification actions in our evolution framework.
-Specifically, we categorize all actions into three main types: Add, Modify, and Remove.
-Each of these categories encompasses a set of fine-grained operations applied to either nodes or structural components (e.g., conditional, loop, parallel).
-In this study, we conducted ablation experiments to evaluate the impact of different operations on the effectiveness of the workflow.
-Specifically, we selected three key operations from the set of possible actions: adding a tool node, modifying a node prompt, and removing a node.
-Firstly, the operation of adding a tool node is intended to acquire information about similar images.
-To assess the necessity of this operation, we removed it from the search space and observed the performance changes.
-Similarly, we applied the same treatment to the operations of modifying a node prompt and removing a node, excluding them from the search space to analyze each operation’s contribution to the overall system performance and effectiveness.
-These experiments will aid in understanding the role and significance of each operation in workflow optimization.
-To assess their individual contribution, we disable one specific action at a time and examine the resulting Top-1 diagnostic accuracy on the Skin Concepts dataset.
-The results are summarized in Table 2, which reveals the relative importance of each operation in enabling effective performance improvements.
-On the one hand, notably, actions related to adding tool node and prompt modification exhibit a particularly strong impact on performance, underscoring their significance in workflow adaptability and optimization.
-On the other hand, we can see that removing nodes does not significantly impact accuracy.
-However, in practice, having too many nodes with similar functions can lead to decreased execution efficiency of the workflow.
+We further evaluate the framework under three different integration levels: node-level (N), node + structural-level (N+S), and full-level (N+S+F).
+The results show that incorporating structural-level modifications leads to performance improvement, with the full-level framework achieving the best results.
+This confirms the effectiveness of hierarchical integration in supporting comprehensive workflow evolution.
+Table 4: Ablation results for integrating different operation levels (N, N+S, N+S+F). The fully-leveled framework delivers the best performance.
+Setting Top-1 Acc. (%) Top-3 Acc. (%) Top-5 Acc. (%) N 26.13 36.94 45.05 N+S 27.03 38.74 46.40 N+S+F 29.28 40.09 50.45
+Together, these experiments provide a comprehensive understanding of how individual operations and integration levels contribute to performance, offering interpretable evidence for the design in our framework.
 5 Conclusion
-Current dermatology-focused multi-agent systems are hindered by rigid, manually designed workflow frameworks that lack the adaptability to meet evolving diagnostic demands and diverse requirements.
-To overcome this limitation, we propose a node-level evolution framework that introduces flexibility through autonomous exploration and development of feasible workflows.
-By enabling dynamic evolution at the node level, our framework allows workflows to adapt organically, facilitating efficient responses to changing conditions and complex diagnostic scenarios.
-This approach significantly enhances flexibility by supporting modular updates, structural adjustments, and parallel processing while autonomously optimizing workflow paths.
-As a result, our framework offers a scalable and adaptive solution for improving diagnostic accuracy, operational efficiency, and responsiveness in dermatological multi-agent systems.
+Current medical diagnostic systems are often constrained by rigid, manually designed workflows that lack the flexibility to adapt to evolving clinical demands and diverse diagnostic requirements.
+To address this limitation, we propose a workflow evolution framework that enables the autonomous exploration and iterative refinement of agent-based diagnostic workflows.
+To enable a more responsive and robust diagnostic workflow, our framework introduces dynamic adaptability at both the structural and prompt levels, allowing it to evolve in response to diagnostic errors.
+In summary, our framework provides a scalable and adaptive solution for advancing multi-agent systems in diagnostic tasks.
+It not only bridges the gap between static, expert-defined architectures and dynamic, self-improving methods, but also lays the groundwork for future adaptive diagnostic systems.